### v1

# Machine-learned RG-improved gauge actions and classically perfect gradient flows

Abstract
Extracting continuum properties of quantum field theories from discretized spacetime is challenging due to lattice artifacts. Renormalization-group (RG)-improved lattice actions can preserve continuum properties, but are in general difficult to parameterize. Machine learning (ML) with gauge-equivariant convolutional neural networks provides a way to efficiently describe such actions. We test a machine-learned RG-improved lattice gauge action, the classically perfect fixed-point (FP) action, for four-dimensional SU(3) gauge theory through Monte Carlo simulations. We establish that the gradient flow of the FP action is free of tree-level discretization effects to all orders in the lattice spacing, making it classically perfect. This allows us to test the quality of improvement of the FP action, without introducing additional artifacts. We find that discretization effects in gradient-flow observables are highly suppressed and less than 1% up to lattice spacings of 0.14 fm, allowing continuum physics to be extracted from coarse lattices. The quality of improvement achieved motivates the use of the FP action in future gauge theory studies. The advantages of ML-based parameterizations also highlight the possibility of realizing quantum perfect actions in lattice gauge theory.
April 22, 2025
Introduction — The Standard Model of particle physics gives excellent agreement between experimental measurements and theoretical predictions to very high precision. A central element is Quantum Chromodynamics (QCD), the quantum field theory of quark-gluon interactions, where Monte Carlo (MC) simulations play a crucial role in testing nonperturbative dynamics [1]. A necessary step in lattice simulations is spacetime discretization, which ultimately must be removed to obtain physical continuum predictions. However, the extrapolation to the continuum limit is often the dominant source of uncertainty in lattice studies and remains a serious challenge: reducing the lattice spacing $a\rightarrow 0$, simulations become less efficient [2], requiring longer Markov chains for high statistical accuracy (critical slowing down) and to fully explore the phase space (topological freezing). These systematic effects are a barrier to increasing the precision of lattice predictions for the Standard Model. To overcome this hurdle, one can use improved lattice actions [3, 4, 5], designed to remove the leading lattice artifacts, often through a perturbative expansion in $a$. The efficacy of the improvement must be empirically tested via MC simulations. The problem has also stimulated the study of various alternatives, such as normalizing flows, which use machine learning (ML) to transform random samples to target distributions [6, 7, 8, 9, 10, 11, 12, 13, 14]. However, normalizing flows face significant challenges, such as achieving high efficiency in four-dimensional quantum field theories—measured by effective sample size—and ensuring that models trained at one lattice spacing can generalize effectively to generate ensembles at other lattice spacings. Combined with the persistent difficulties of extrapolating to the continuum limit, these challenges underscore the urgent need for new methods that can achieve accurate results with small lattice artifacts already on coarse lattices, thus evading critical slowing down and topological freezing.
In this Letter, we employ an alternative approach, based on fixed-point (FP) lattice actions [15] which are implicitly and nonperturbatively defined using the properties of Wilson’s renormalization group (RG) for asymptotically free theories [16, 17, 18]. It is used to extract continuum quantities from coarse lattices with significantly reduced lattice artifacts. We demonstrate the effectiveness of the FP action in MC simulations of four-dimensional SU(3) gauge-field theory, parameterized using a convolutional neural network with exact gauge invariance [19]. This result is made possible by our new theoretical insight, presented here, that FP actions yield a classically perfect gradient flow, i.e., without tree-level lattice artifacts to all orders in the lattice spacing. This result also builds on the success of our recent work [20, 21, 22], where we demonstrated that ML-based parameterizations can accurately capture the FP action across a wide range of gauge-field fluctuations—from very fine to very coarse lattice spacings. The resulting parameterization achieves an average relative error in the action value of less than $0.2\%$, representing an improvement by over an order of magnitude compared to previous approaches developed earlier [23, 24, 25]. The ML framework also enables accurate parameterization of the derivatives of the FP action w.r.t. the gauge fields, which are crucial for both efficient network training and the practical implementations of MC algorithms. The FP action can be generally used in any lattice-QCD study and provides a new way to reduce discretization artifacts on coarse lattices, making it broadly useful.
Renormalization group and fixed-point actions — For SU($N$) gauge theories, an RG transformation of a lattice action $\mathcal{A}$ from a fine to a coarse lattice can be defined through
$\exp(-\beta^{\prime}\mathcal{A}^{\prime}[V])=\int\mathcal{D}U\exp(-\beta\{\mathcal{A}[U]+\mathcal{T}[U,V]\}),$ (1)
where $\beta=2N/g^{2}$ and the fine lattice configuration $U$ is connected to the coarse configuration $V$ through the RG blocking kernel
$\mathcal{T}[U,V]=-\kappa\sum_{n_{B},\mu}\Bigl{\{}{\rm ReTr}(V_{n_{B},\mu}Q_{n_{B},\mu}^{\dagger}[U])-\mathcal{N}^{\infty}_{\mu}\Bigr{\}}.$ (2)
The variables $Q_{n_{B},\mu}$ are blocked links constructed from the fine links $U_{n,\mu}$ and the normalization constant $\mathcal{N}^{\infty}_{\mu}=\max_{W\in\text{SU}(N)}\left[{\rm ReTr}(WQ_{n_{B},\mu}^{\dagger})\right]$ guarantees the invariance of the partition function after an RG blocking step. The blocking details and parameter $\kappa$ can be freely chosen to maximize the compactness of the FP action [24] and in this work we follow the RG blocking therein.
In the limit $1/g^{2}\rightarrow\infty$ of asymptotically free quantum field theories, the FP of the RG is scale-invariant and defines the continuum limit of the gauge theory. Moreover, the path integral on the RHS of Eq. (1) reduces to a saddle point problem yielding the classical FP equation [15]
$\mathcal{A}^{\rm FP}[V]=\min_{\{U\}}\left[\mathcal{A}^{\rm FP}[U]+\mathcal{T}[U,V]\right].$ (3)
The resulting FP action $\mathcal{A}^{\rm FP}$ is classically perfect, i.e., it has no lattice artifacts at tree level and exactly reproduces continuum classical properties at finite lattice spacing, e.g., dispersion relation [15], topology [23, 26], or index theorem [27], chirality [28, 29] when fermions are included. A key theoretical result of this work is the derivation that the gradient flow of the FP action is also classically perfect.
Classically perfect gradient flow — To rigorously test the quality of improvement in any lattice action, it is essential to use precisely measurable observables that avoid introducing significant cutoff effects. Observables based on the gradient flow (GF) of gauge fields [30] are a natural choice, given their small systematics and high statistical accuracy. In addition, due to the renormalization properties of the GF [31], these observables are widely used for the setting of physical scales and the measurement of the $\beta$-function of the gauge coupling. We show here that the GF based on the FP action has the additional desirable property of being classically perfect: it has no tree-level discretization effects to any order in the lattice spacing $a$.
[…]
$\frac{dA_{\mu}(t)}{dt}=-\frac{\delta{\cal A}}{\delta A_{\mu}(t)},$ (4)
[…]
$t^{2}\langle E(t)\rangle=\frac{3(N^{2}-1)g^{2}}{128\pi^{2}}\left(1+{\cal O}(g^{2})\right),$ (5)
[…]
$t^{2}\langle E(t)\rangle=\frac{3(N^{2}-1)g_{0}^{2}}{128\pi^{2}}\left(C(a^{2}/t)+{\cal O}(g_{0}^{2})\right),$ (6)
where
$C(a^{2}/t)=\frac{64\pi^{2}t^{2}}{3}\times\\ \int_{-\pi/a}^{\pi/a}\frac{d^{4}p}{(2\pi)^{4}}{\rm Tr}\left[e^{-t({\cal S}^{f}+{\cal G})}({\cal S}^{g}+{\cal G})^{-1}e^{-t({\cal S}^{f}+{\cal G})}{\cal S}^{e}\right]$ (7)
contains the tree-level lattice artifacts as deviations from 1. Here, the action density in momentum space to quadratic order in the fields is $A_{\mu}(p){\cal S}_{\mu\nu}(p)A_{\nu}(-p)$, and ${\cal S}^{e},{\cal S}^{f},{\cal S}^{g}$ correspond to the (possibly different) choices of the discretized action for the action-density observable $E$, for the GF, and for the MC simulation, respectively, while ${\cal G}$ is a gauge fixing term. For any choice of action discretizations, the value of $C(a^{2}/t)$ can be calculated and used to define a tree-level improved coupling at finite lattice spacing [34] (see also [35]). Using the same lattice action for all three gives
$C(a^{2}/t)=\frac{64\pi^{2}t^{2}}{3}\int_{-\pi/a}^{\pi/a}\frac{d^{4}p}{(2\pi)^{4}}{\rm Tr}\left[e^{-2t({\cal S}+{\cal G})}\right],$ (8)
since the various factors under the trace commute with each other.
[…]
$C(a^{2}/t)=\frac{64\pi^{2}t^{2}}{3}\cdot 3\left(\int_{-\infty}^{+\infty}\frac{dp}{2\pi}e^{-2tp^{2}}\right)^{4}=\frac{64\pi^{2}t^{2}}{(\sqrt{8\pi t})^{4}}=1,$ (9)
i.e., the exact FP action has no tree-level cutoff dependence in $t^{2}\langle E(t)\rangle$; lattice artifacts appear in loop corrections, i.e., as quantum effects, at ${\cal O}(a^{2}g_{0}^{2})$. This is a new example of FP actions keeping continuum classical properties intact at finite lattice spacing.
Machine learning architecture — The practical limitations in using the FP action in MC simulations arise from two factors: the size of quantum lattice artifacts, and the inherent imperfections in the fixed-point parameterization for four-dimensional SU(3) gauge theory. The latter is addressed here using a gauge-equivariant convolutional neural network [19] and ML techniques. The ML architecture designed for the FP action is described in detail in [21]. In brief, the main building block of the neural network is a bilinear convolution which computes various products of Wilson loops, starting from $1\times 1$ plaquettes and the links $U_{n,\mu}$ of a gauge-field configuration as input. Gauge covariance is exactly maintained throughout by accounting for parallel transport in the convolution. As the network depth increases, larger and more complex closed loops are generated. The final layer is a trace giving the action value as output. The network respects translational symmetry and is (ultra-)local, which allows it to be used on any lattice size. Additionally, our architecture is set up to reproduce a perturbatively defined FP action for smooth fields [24] in order to guarantee the correct continuum limit.
The goal of training a model is to approximate the action values according to the FP equation in Eq.(3). A crucial element is that the exact derivatives of ${\cal A}^{\rm FP}$ with respect to gauge links $U_{n,\mu}$ can be easily obtained from the FP equation, with each gauge configuration giving 8×4×8$\times$ 4$\times$8 × 4 ×(volume) data for learning. In the network, the derivatives can be calculated exactly via backpropagation. The loss function for training hence contains both action values and derivatives. The best architecture has been obtained from a hyperparameter scan and consists of three gauge-covariant convolutional layers. It has been trained on equilibrated configurations for a variety of lattice sizes and a large range of coupling parameters from very smooth to coarse gauge-field fluctuations, and has further been improved by finetuning on instanton configurations. Our best model reaches an average relative error in the action value of less than 0.2% with the exact derivatives more accurately reproduced than in previous parameterizations [25] across all lattice spacings. The fidelity of the model for very smooth configurations is particularly important for the GF because the gauge-field fluctuations become very smooth at large flow time.
Fixed-point action simulations — The availability of accurate gauge-field derivatives from our ML parameterization model makes the Hybrid Monte Carlo (HMC) algorithm a suitable choice for simulations, where artificial conjugate momenta are introduced to update the gauge links through molecular dynamics [36, 37]. The simulations can be accelerated with a coarser choice of time step when solving the equations of motion [38] using in particular the 4MN4FP variant of Omelyan integration scheme [39]. The GF for gauge links $dU_{\mu}/dt=-i(\delta{\cal A}^{\rm FP}/\delta U_{\mu})U_{\mu}$ is solved using the third-order Runge-Kutta integrator [30]. The ML network is memory intensive; the largest lattice volumes we simulate are $18^{4}$, for which we can generate a new HMC trajectory in $\sim 4$ minutes on one NVidia A100 SXM6 64GB GPU. With the FP action, a smaller lattice volume is compensated by coarser lattice spacing to yield near-continuum-physics results in large enough physical volume. As a consistency check of continuum predictions with the FP action, we also simulate four-dimensional SU(3) gauge theory using either Wilson or tree-level Symanzik improved gauge actions, measuring with Wilson GF, using the publicly available openqcd package [40].
Scaling tests — The GF provides several physical quantities that can be extracted. For example, a reference scale $t_{c}/a^{2}$ can be set through the condition $t_{c}^{2}\langle E(t_{c})\rangle=c$, where a common choice is $c=0.3$ [30]. An alternative scale $w_{d}/a$ is set through the derivative condition
$t\cdot\frac{d}{dt}\left[t^{2}\langle E(t)\rangle\right]_{t=t_{d}}=d,\hskip 14.22636ptw_{d}=\sqrt{t_{d}},$ (10)
where $d=0.3$ is also frequently used [41]. At finite lattice spacing, the values of $t_{c}/a^{2}$ and $w_{d}/a$ are determined by the choice of bare coupling $6/g_{0}^{2}$ at which MC simulations are performed, as well as the action discretization used in the GF and the action density. Ratios of scales, e.g., $t_{c}/w_{d}^{2}$ or $t_{d}/t_{c}$, are physical properties, which have a well-defined value in the continuum limit independent of the discretizations employed. The necessary consistency in the continuum limit is a consequence of universality and provides a stringent test of the FP simulation results.
Figure 1: Continuum-limit extrapolations for the ratios $t_{0.3}/w_{0.3}^{2}$ and $t_{0.5}/t_{0.3}$. Results from Wilson and Symanzik MC simulations are shown using plaquette and clover discretizations of the action density.
In addition, the renormalizability of the GF action density $E(t)$ allows one to define a scheme for a renormalized coupling. For SU(3) gauge theory the GF scheme is defined by
$t^{2}\langle E(t)\rangle\doteq\frac{3g_{\rm GF}^{2}(t)}{16\pi^{2}},$ (11)
where the GF renormalized coupling $g^{2}_{\rm GF}(t)$ runs with the RG scale $\mu=1/\sqrt{8t}$. The $\beta$-function of the gauge coupling $\mu^{2}\cdot dg^{2}_{\rm GF}/d(\mu^{2})=-t\cdot dg^{2}_{\rm GF}/dt$ can be measured directly from the GF at the desired value of the coupling [42, 43]. For example, the choice $c=0.3$ corresponds to strong coupling $g^{2}_{\rm GF}\approx 15.79$. In the perturbative regime, it can be compared with the known 3-loop calculation [44].
Repeating the procedure for a set of gauge ensembles at different bare couplings, the continuum limit is taken as $a^{2}/t_{c}\rightarrow 0.$ We present in Figures 1 and 2 results for the ratios $t_{0.3}/w_{0.3}^{2}$ and $t_{0.5}/t_{0.3}$, and the $\beta$-function $t\cdot dg_{\rm GF}^{2}/dt$ at fixed $g^{2}_{\rm GF}\approx 15.79$, using FP, Wilson and Symanzik action simulations. Further examples can be found in the Supplemental Material. The leading lattice artifacts are expected to be ${\cal O}(a^{2})$ for the Wilson data, ${\cal O}(a^{4})$ for the Symanzik data, and ${\cal O}(g^{2}a^{2})$ for the FP data, and we show various possible fits (polynomial in $a^{2}/t_{0.3}$) describing the continuum-limit extrapolations. [^1] We note that the leading ${\cal O}(a^{4})$ behaviour of the Symanzik data is masked by the ${\cal O}(a^{2})$ discretization effects from the flow action (Wilson) and the action density (plaquette and clover). The fit curves are shaded according to their Akaike Information Criterion (AIC) weight [45], cf. Supplemental Material.
The FP data for $t_{0.3}/w_{0.3}^{2}$ and the $\beta$-function show mild variation with lattice spacing, with artifacts smaller than $1\%$ at $a\simeq 0.14$ fm, using $\sqrt{t_{0.3}}=0.1679\,\mathrm{fm}$ [46] to convert from lattice to physical scales. The ratio $t_{0.5}/t_{0.3}$ has essentially no lattice-spacing dependence below $a\lesssim 0.14$ fm using the FP action. By comparison, discretization errors with the Wilson or Symanzik action and either plaquette or clover version of the action density $E$ are much larger, hence finer lattice spacings and larger lattice volumes are necessary to reach a scaling window of ${\cal O}(a^{2})$ behavior while simultaneously keeping finite-volume effects small.
Figure 2: The $\beta$-function of the GF renormalized coupling at $g_{\rm GF}^{2}=15.79$ with highly suppressed lattice artifacts in the results for the FP compared to the Wilson and Symanzik gauge actions.
To estimate systematic effects in the continuum extrapolation, such as choosing the lattice-spacing range over which to fit or the order of the fit polynomial in $a^{2}/t_{0.3}$, we use a procedure based on the AIC, cf. Supplemental Material.
The advantageous features of the FP action are visible in other choices of $c$ and $d$, cf. Supplemental Material for further examples. In each case, FP lattice artifacts are small, and in absolute value the FP continuum predictions are accurate to per mille level. We summarize a comparison of our continuum results using the FP, Wilson or Symanzik action in Figure 3, for values of $c$ and $d$ employed by other groups [47, 48, 49, 50, 51, 52, 53, 54]. There is clear consistency between FP, Wilson and Symanzik results. Simulating on smaller coarser lattices, one can measure precisely the continuum properties of the gauge theory with the FP approach.
Conclusions and outlook — The explicit form of the exact RG-improved FP action is unknown and must be parameterized as a solution of the nonperturbative FP equation. Our previous study [21] demonstrated that ML networks can describe the FP action for four-dimensional SU(3) gauge theory to much higher accuracy than previously possible [55, 24, 25]. In this work, we show that the FP action defines a classically perfect gradient flow, i.e., without tree-level lattice artifacts to all orders in the lattice spacing. As a consequence, the FP action successfully reduces lattice artifacts in gradient-flow observables, paving the way for various future applications. As an RG-improved lattice action, it can be used in any lattice-QCD study. One application is a new measurement of the $\Lambda$-parameter of SU(3) pure gauge theory, as a consistency check of recent results [56, 52, 51] which differ from the older literature average [1], and as quantitative input for the determination of the strong coupling $\alpha_{S}(M_{Z})$ through the decoupling of quarks [57]. Another possibility is using ML to parameterize the FP Dirac operator [28, 58, 59] and reduce cutoff effects in lattice-QCD simulations with dynamical quarks. A third is using ML to construct quantum perfect actions [15], where all lattice artifacts are removed. As a general lesson, our study shows a new way in which ML can enhance particle physics simulations.
Figure 3: The comparison of continuum predictions for four-dimensional SU(3) gauge theory from MC simulations using either the FP, Wilson or tree-level Symanzik improved lattice action shows very good consistency. The $\beta$-function results are rescaled by a factor of 50 for visibility.
I Acknowledgements
Acknowledgements.
K.H. wishes to thank the AEC and ITP at the University of Bern for their support. U.W. acknowledges funding from the Swiss National Science Foundation (SNSF) project No. 200020_208222. A.I. and D.M. acknowledge funding from the Austrian Science Fund (FWF) projects P 32446, P 34455, P 34764 and PAT3667424. K.H. acknowledges funding from the US National Science Foundation under grants 2014150 and 2412320. The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and LEONARDO at CINECA, Italy, via an AURELEO (Austrian Users at LEONARDO supercomputer) project, and UBELIX (http://www.id.unibe.ch/hpc), the HPC cluster at the University of Bern.